# The Riemann Hypothesis is false

### v40

@@
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$. Thus by Abel summation it follows that
$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq\varepsilon m-1}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (35)

### v41

@@
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant, $\chi(s)=\frac{\zeta(s)}{\zeta(1-s)}$ and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1+it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$. Infact, by adapting the argument between (18) and (34), one finds that $\sum_{0<r\leq J}e^{2\pi if(r)}=o(1)$ uniformly for $11\leq J\leq b=m\varepsilon-$ 1. Thus by Abel summation it follows that
$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq b}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (35)